OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 3) and a radius of 1. x2 + (y – 3)2 = 1 x2 + (y – 3)2 = 2 x2 + (y + 3)2 = 1 x2 + (y + 3)2 = 2

OpenStudy (anonymous):

Note that equation of the circle is as such\[(x-a)^2+(y-b)^2 = r^2\]

OpenStudy (anonymous):

Where your center is (a,b)

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

Not quite, look at the radius

OpenStudy (anonymous):

x2 + (y – 3)2 = 1?

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

thanks